# On $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ and Frege Systems

### v13

@@ Abstract
We prove in this paper that there is a language $L_{d}$ accepted by some nondeterministic Turing machines but not by any ${\rm co}\mathcal{NP}$-machines (defined later). Then we further show that $L_{d}$ is in $\mathcal{NP}$, thus proving that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$. The main techniques used in this paper are lazy-diagonalization and the novel new technique developed in the author’s recent work [Lin21]. Further, since there exists some oracle $A$ such that $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, we then explore what mystery lies behind it and show that if $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, then the set of all polynomial-time co-nondeterministic oracle Turing machines with oracle $A$ is not enumerable, thus showing that the technique of lazy-diagonalization is not applicable for the first half of the whole step to separate $\mathcal{NP}^{A}$ from ${\rm co}\mathcal{NP}^{A}$. As a by-product, we reach the important result that $\mathcal{P}\neq\mathcal{NP}$ [Lin21] once again, which is clear from the above outcome and the well-known fact that $\mathcal{P}={\rm co}\mathcal{P}$. Next, we show that the complexity class ${\rm co}\mathcal{NP}$ has intermediate languages, i.e., there exists a language $L_{inter}\in{\rm co}\mathcal{NP}$, which is not in $\mathcal{P}$ and not ${\rm co}\mathcal{NP}$-complete. We also summarize other direct consequences implied by our main outcome, such as $\mathcal{NEXP}\neq{\rm co}\mathcal{NEXP}$, $\mathcal{BPP}\neq\mathcal{NEXP}$ and there exists no super proof system. Lastly, we show a lower bounds result for Frege proof systems, i.e., no Frege proof systems can be polynomially bounded.
@@ 8.1 Lower Bounds for Frege Systems
Initially, propositional proof complexity has been primarily concerned with proving lower bounds (even conditional) for the length of proofs in propositional proof systems, which is extremely interesting and well justified in its own right, with the ultimate goal of settling whether $\mathcal{NP}={\rm co}\mathcal{NP}$ (see e.g. [CR79]). In fact, as mentioned in Section 1, such research directions are called Cook’s Program for separating $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ (see e.g. [Bus12]): Prove superpolynomial lower bounds for proof lengths in stronger and stronger propositional proof systems until they are established for all abstract proof systems. However, our approach is “to do the opposite”, that is, we first prove that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, and then apply it to prove the lower bounds on the length of proofs in Frege proof systems.
@@ 9 Conclusions
thus resolving a very important and long-standing conjecture in computational complexity theory. Specifically, we first showed that the set of all ${\rm co}\mathcal{NP}$-machines is enumerable. Then, we constructed a universal nondeterministic Turing machine $D$ to simulate all ${\rm co}\mathcal{NP}$-machines and to do the lazy-diagonalization operation to create a language $L_{d}$ which is not in ${\rm co}\mathcal{NP}$. Next, we applied the technique developed in the author’s recent work [Lin21] to show that the language $L_{d}$ accepted by the universal nondeterministic Turing machine $D$ is in fact in $\mathcal{NP}$, thus establishing the desired main conclusion of this paper.
@@
which suggests that (in the light of the point of view given in [BGS75]) ordinary diagonalization techniques (lazy-diagonalization) are not capable of proving $\mathcal{NP}\neq{\rm co}\mathcal{NP}$; see [BGS75]. We explored the mystery behind this phenomenon by proving that if $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, then the set of all ${\rm co}\mathcal{NP}$-machines with oracle $A$ is in fact not enumerable, thus showing that the ordinary diagonalization techniques (i.e., lazy-diagonalization) will generally not apply to the relativized versions of the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem.
As a by-product, the important result that $\mathcal{P}\neq\mathcal{NP}$ which is shown in [Lin21], also follows. The above outcome can be seen as, in a sense, the direct or indirect consequence of the proof procedures of
@@
Note again that by Corollary 3 we in fact have shown $\mathcal{P}\neq{\rm co}\mathcal{NP}$ because if $\mathcal{P}={\rm co}\mathcal{NP}$, then $\mathcal{P}={\rm co}\mathcal{P}=\mathcal{NP}$, a contradiction to Corollary 3. As a matter of fact and interestingly, we have also shown by Theorem 6 that there exists a ${\rm co}\mathcal{NP}$-intermediate language in ${\rm co}\mathcal{NP}$, which is not in $\mathcal{P}$ nor ${\rm co}\mathcal{NP}$-complete, i.e., the complexity class ${\rm co}\mathcal{NP}$ also has a rich structure. We have also obtained other interesting consequences, such as Corollary 4 (i.e., $\mathcal{NEXP}\neq{\rm co}\mathcal{NEXP}$), Corollary 5 (i.e., $\mathcal{BPP}\neq\mathcal{NEXP}$) and Corollary 7, which states that there is no super proof system.
As mentioned in the introduction section, the problem of $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ closely connects with the field of proof complexity. In particular, we have shown that no Frege proof systems are polynomially bounded (Theorem 8), thus answering an open problem in [Pud08]. As pointed out in Remark 8.2, we do the opposite direction of Cook’s program. That is, we first prove that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, and then apply it to prove the lower bounds on the length of proofs in Frege proof systems. Note that one of the key points to show the Theorem 8 is that nondeterministic Turing machines can simulate Frege proofs (see [Bus99]). But we do not know whether the nondeterministic Turing machines can simulate Extended Frege proofs. For more details about Extended Frege Systems, please consult the references [CR79, CN10, Kra95].
Lastly, there are many other interesting open problems in the field of proof complexity that we did not touch on in this paper, and the interested reader is referred to these references [Bus12, CR74, CR79, CN10, Kra95, Kra19, Pud08, Raz03, Raz15] for those excellent open problems.

### v14

@@ Abstract
We prove in this paper that there is a language $L_{d}$ accepted by some nondeterministic Turing machines but not by any ${\rm co}\mathcal{NP}$ machines. Then we further show that $L_{d}$ is in $\mathcal{NP}$, thus proving that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$. The main techniques used in this paper are lazy diagonalization and the novel new technique developed in the author’s recent work [Lin21]. Further, since there exists some oracle $A$ such that $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, we then explore what mystery lies behind it and show that if $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, then the set of all ${\rm co}\mathcal{NP}^{A}$ machines with oracle $A$ is not enumerable, thus showing that the technique of lazy diagonalization is not applicable for the first half of the whole step to separate $\mathcal{NP}^{A}$ from ${\rm co}\mathcal{NP}^{A}$. Our main result has profound implications, such as $\mathcal{P}\neq\mathcal{NP}$ and a lower bounds result for Frege proof systems (i.e., no Frege proof systems can be polynomially bounded), etc.
@@ 8.1 Lower Bounds for Frege Systems
Initially, propositional proof complexity has been primarily concerned with proving lower bounds (even conditional) for the length of proofs in propositional proof systems, which is extremely interesting and well justified in its own right, with the ultimate goal of settling whether $\mathcal{NP}={\rm co}\mathcal{NP}$ (see e.g. [CR79]). In fact, as mentioned in Section 1, such research directions are called Cook’s Program for separating $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ (see e.g. [Bus12]): Prove superpolynomial lower bounds for proof lengths in stronger and stronger propositional proof systems until they are established for all abstract proof systems. However, our approach is “to do the opposite,” that is, we first prove that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, and then apply it to prove the lower bounds on the length of proofs in Frege proof systems.
@@ 9 Conclusions
thus resolving a very important and long-standing conjecture in computational complexity theory and in proof complexity theory. Specifically, we first showed that the set of all ${\rm co}\mathcal{NP}$ machines is enumerable. Then, we constructed a universal nondeterministic Turing machine $D$ to simulate all ${\rm co}\mathcal{NP}$ machines and to do the lazy-diagonalization operation to create a language $L_{d}$ which is not in ${\rm co}\mathcal{NP}$. Next, we applied the technique developed in the author’s recent work [Lin21] to show that the language $L_{d}$ accepted by the universal nondeterministic Turing machine $D$ is in fact in $\mathcal{NP}$, thus establishing the desired main conclusion of this paper.
@@
which suggests that (in light of the point of view given in [BGS75]) ordinary diagonalization techniques (lazy-diagonalization) are not capable of proving $\mathcal{NP}\neq{\rm co}\mathcal{NP}$; see [BGS75]. We explored the mystery behind this phenomenon by proving that if $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, then the set of all ${\rm co}\mathcal{NP}$ machines with oracle $A$ is in fact not enumerable, thus showing that the ordinary diagonalization techniques (i.e., lazy-diagonalization) will generally not apply to obtaining the lower bounds result and further to the relativized versions of the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem.
As a by-product, the important result that $\mathcal{P}\neq\mathcal{NP}$, which is shown in [Lin21], also follows. The above outcome can be seen as, in a sense, the direct or indirect consequence of the proof procedures of
@@
Note again that by Corollary 3 we in fact have shown $\mathcal{P}\neq{\rm co}\mathcal{NP}$ because if $\mathcal{P}={\rm co}\mathcal{NP}$, then $\mathcal{P}={\rm co}\mathcal{P}=\mathcal{NP}$, a contradiction to Corollary 3. As a matter of fact and interestingly, we have also shown by Theorem 6 that there exists a ${\rm co}\mathcal{NP}$-intermediate language in ${\rm co}\mathcal{NP}$, which is not in $\mathcal{P}$ nor ${\rm co}\mathcal{NP}$-complete, i.e., the complexity class ${\rm co}\mathcal{NP}$ also has a rich structure. We have also obtained other interesting consequences, such as Corollary 4 (i.e., $\mathcal{NEXP}\neq{\rm co}\mathcal{NEXP}$), Corollary 5 (i.e., $\mathcal{BPP}\neq\mathcal{NEXP}$), and Corollary 7, which states that there is no super proof system.
As mentioned in the introduction section, the problem of $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ closely connects with the field of proof complexity. In particular, we have shown that no Frege proof systems are polynomially bounded (Theorem 8), thus answering an open problem in [Pud08]. As pointed out in Remark 8.2, we do the opposite direction of Cook’s program. That is, we first prove that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$ and then apply it to prove the lower bounds on the length of proofs in Frege proof systems. Note that one of the key points to show the Theorem 8 is that nondeterministic Turing machines can simulate Frege proofs (see [Bus99]). But we do not know whether the nondeterministic Turing machines can simulate extended Frege proofs. For more details about Extended Frege Systems, please consult the references [CR79, CN10, Kra95].
In addition, there are many other interesting open problems in the field of proof complexity that we did not touch on in this paper, and the interested reader is referred to these references [Bus12, CR74, CR79, CN10, Kra95, Kra19, Pud08, Raz03, Raz15] for those excellent open problems.